# Evaluating LLMs on Entity Disambiguation in Tables

### v2

@@ 3. Selected Approaches
Table 1 lists the datasets used to train the selected approaches; in particular, for TableLlama we use the published model on HugginFace, while we train TURL and Alligator on the same datasets used in the original paper. We use the term ”pre-train” to refer to this main training phase to distinguish it from subsequent fine-tuning (cfr. Section 4.4). Dagobah is not listed because is not based on training.
@@ 5.3. Discussion
Generative GTUM approaches with a high number of parameters have interesting properties in terms of accuracy, generalisation, and robustness to the number of candidates that are processed and the table metadata, especially on MOOD and OOD data, as found in Sections 5.1 and 5.2; however, specific STI approaches trained on in-domain data, and optionally fine-tuned on moderately-out-of-domain data, using a tiny fraction of these parameters can still be on par with generalistic approaches while being faster and occupying less memory. Generative approaches like GPT-4o and TableLlama are more promising than encoder-based ones in terms of performance, with a negligible risk of hallucinations; however, the comparison considered models of uncomparable size (7B vs 300M for TableLlama and TURL resp.), with TURL being the fastest approach among those that were tested. Regarding scalability, long context allows the encapsulation of large tables with a high number of candidates, but some tables cannot fit in the context, and when they do, the computation time and occupied memory increase proportionally to the table size: considering a small context table (few rows above and below the row containing the mention to be linked) to be fed to the model could be effective. On the budget side, training models of the size of TableLlama has still enormous costs (48 A100-80GB for 9 training days), so devising generative methods based on smaller LLMs, e.g., Phi (abdin2024phi3, ), could be an interesting research direction, although more sophisticated approaches may be needed to achieve the same level of generalisation and reliability. Moreover, Figure 6 shows the accuracy achieved by the tested approaches, both before and after MOOD fine-tuning, with respect to the overall price [^12] For TableLlama, TURL and Alligator, prices are estimated using the Azure Pricing Calculator, selecting the NC24ads_a100_v4 for running GPU inference with TableLlama and TURL, and D48ds_v5 for running CPU inference and hosting Alligator. GPT-4o and GPT-4o-mini prices are estimated from the OpenAI prices page, considering a token as being equal to 4 characters. paid to host machines to run inference on all the test data. The figure shows how GPT-4o, while being the best-performing model on average, is also the most expensive one; GPT-4o-mini, TableLlama, TableLlama-FT, and Alligator-FT are comparable in terms of accuracy and price paid. These results suggest that generative GTUM approaches are best suited when the tables to be annotated contain rich, highly semantic textual data. GPT-4o is the most accurate among these methods but also the most expensive. GPT-4o-mini and TableLlama offer a fair compromise between accuracy, cost, and computational demand. On the other hand, Alligator is more appropriate when signals from the knowledge graph (KG) have a significant impact on the final accuracy, especially when time and budget are limited. When fine-tuned, Alligator performs comparably to GPT-4o-mini and TableLlama in terms of accuracy and cost.
@@
Figure 6. Per-approach accuracy over price. The price is computed given the overall time spent (in minutes) to run the evaluation over all the test data (15205 mentions).

### v3

@@ 3. Selected Approaches
Table 1 lists the datasets used to pre-train the selected approaches; in particular, for TableLlama we use the published model on HugginFace, while we train TURL and Alligator on the same datasets used in the original paper. We use the term ”pre-train” to refer to this main training phase to distinguish it from subsequent fine-tuning (cfr. Section 4.4). Dagobah is not listed because is not based on training.
@@ 4.4. Training and implementation details
Evaluation and inference. Following our primary objective, all selected approaches were evaluated on the same test data (refer to Table 3). Both TableLlama and TURL were run with a batch size of 1, though this setting has distinct interpretations: for TURL, it means processing one table at a time, while for TableLlama, it means processing a single mention at a time. The GPT-4o and GPT-4o-mini models were also run with TableLlama-style prompts, meaning they process one mention at a time. In contrast, Alligator was run with its default configuration, using 8 parallel processes, each handling up to 50 rows per table. Lastly, Dagobah was executed with its default hyperparameters, as specified in its official GitHub repository.
@@ 5.3. Discussion
Generative GTUM approaches with a high number of parameters have interesting properties in terms of accuracy, generalisation, and robustness to the number of candidates that are processed and the table metadata, especially on MOOD and OOD data, as found in Sections 5.1 and 5.2; however, specific STI approaches trained on in-domain data, and optionally fine-tuned on moderately-out-of-domain data, using a tiny fraction of these parameters can still be on par with generalistic approaches while being faster and occupying less memory. Generative approaches like GPT-4o and TableLlama are more promising than encoder-based ones in terms of performance, with a negligible risk of hallucinations; however, the comparison considered models of uncomparable size (7B vs 300M for TableLlama and TURL resp.), with TURL being the fastest approach among those that were tested. Regarding scalability, long context allows the encapsulation of large tables with a high number of candidates, but some tables cannot fit in the context, and when they do, the computation time and occupied memory increase proportionally to the table size: considering a small context table (few rows above and below the row containing the mention to be linked) to be fed to the model could be effective. On the budget side, training models of the size of TableLlama has still enormous costs (48 A100-80GB for 9 training days), so devising generative methods based on smaller LLMs, e.g., Phi (abdin2024phi3, ), could be an interesting research direction, although more sophisticated approaches may be needed to achieve the same level of generalisation and reliability. Moreover, Figure 6 shows the accuracy achieved by the tested approaches, both before and after MOOD fine-tuning, with respect to the average price [^12] For TableLlama, TURL and Alligator, prices are estimated using the Azure Pricing Calculator, selecting the NC24ads_a100_v4 for running GPU inference with TableLlama and TURL, and D48plds_v5 for running CPU inference and hosting Alligator. GPT-4o and GPT-4o-mini prices are estimated from the OpenAI prices page, considering a token as being equal to 4 characters. paid to host machines to run inference on all the test data. The figure shows how GPT-4o, while being the best-performing model on average, is also the most expensive one; GPT-4o-mini, TableLlama, TableLlama-FT, and Alligator-FT are comparable in terms of accuracy and price paid, although TableLlama costs is slightly higher. These results suggest that generative GTUM approaches are best suited when the tables to be annotated contain rich, highly semantic textual data. GPT-4o is the most accurate among these methods but also the most expensive. GPT-4o-mini offers a fair compromise between accuracy, cost, and computational demand. On the other hand, Alligator is more appropriate when signals from the knowledge graph (KG) have a significant impact on the final accuracy, especially when time and budget are limited. When fine-tuned, Alligator performs comparably to GPT-4o-mini and TableLlama in terms of accuracy with a reduced cost.
@@
Figure 6. Per-approach accuracy over price, on average. The price is computed given the average time spent (in minutes) to run the evaluation over all the test data.
@@ 6. Conclusions and Future Works
7. Acknowledgments
This work has been partially funded by the European innovation action enRichMyData (HE 101070284) and the Italian PRIN project Discount Quality for Responsible Data Science: Human-in- the-Loop for Quality Data (202248FWFS) funded by the European Community - Next Generation EU.